OpenStudy (anonymous):

I will give you a medal if can help me Solve the system using the elimination method. 1/2x+3y=9 1/3x+y=4 Show work PLEASE :)

OpenStudy (anonymous):

muliply the bottom eqaution be -3

OpenStudy (anonymous):

yeah and then add

OpenStudy (anonymous):

you should get: 1/2x+3y=9 +(-1x-3y=-12) --------------- -1/2x=-3

OpenStudy (anonymous):

6

OpenStudy (anonymous):

No, it's 6

OpenStudy (anonymous):

You don't divide by fractions, u multiply by the reciprocal

OpenStudy (anonymous):

y=2

OpenStudy (anonymous):

right, then plug x=6 into in 1/2x+3y=9

OpenStudy (anonymous):

you should get: (1/2)(6)+3y=9

OpenStudy (anonymous):

simplifying: 3+3y=9

OpenStudy (anonymous):

subtract 3 from both sides: 3y=6

OpenStudy (anonymous):

then divide both sides by 3 and get y=2